150 j of heat energy is required to increase the temperature of a body by 6.0°c , calculate the heat capacity of the body and thus it's specific heat capacity if it's mass is 0.1kg

1
Expert's answer
2022-01-30T13:51:36-0500

Answer

Heat required Q=150J

Increment in temperature

"\Delta T= 6°C"

Mass m=0.1Kg

So heat capacity is given by

"C=\frac{ Q}{\Delta T}\\=\frac{ 150}{6}\\=75J/°C"

Specific heat capacity

"C_m=\frac{ Q}{m\Delta T}\\=\frac{ 150}{0.1\times6}\\=150J/°C-Kg"